The SuperB factory

physics prospects and project status

Marcin Chrząszcz

on behave of SuperB Collaboration

Institute of Nuclear Physics PAN Krakow, Poland

put date

Outline

1 Introduction
2 SuperB Infrasctructure

Accelerator Luminosity

3 Detector

SVT DCH DIRC EMC and IFR

4 Physics

Rare B Physics

Introduction

B factories

B factories achived a great success over the dozen years. A natural continuation of this project are Super Flavor Factories. Super Flavor Factories

1 Data 75ab^{−1} . 2 Luminosity 10^{36}cm^{−2} s^{−1} . 3 Flexibility to run on charm threshold with luminosity 10^{35}cm^{−2} s^{−1} . 4 Logitudanal polarization of electron beam 80%. 5 Upgradet Babar detector. 6 Start of data taking: 2018. 7 10ab^{−1} peer year.

SuperB Infrasctructure  Accelerator

SuperB Infrasctructure  Luminosity

Quest for Luminosity

L ∝ \frac{1}{\sqrt{β}_{y}} , Φ ≈ \frac{σ_{z}}{σ_{x}} \frac{θ}{2}

L ∝ \frac{1}{\sqrt{β}_{y}} , Φ ≈ \frac{σ_{z}}{σ_{x}} \frac{θ}{2}

Detector

Recycling

SuperB detector is based on Babar.

Detector  SVT

Silicon Vertex Tracker

Five layers(1-5) of double-sided silicon strip detectors. Radial span 3 − 15 cm. Upgrade the electronics for faster readout. Additional Layer 0:

1 Radius ≈ 1.5cm .

2 Low material budget: X_{0} = 0.5%.

3 Two possible technologies: Hybrid Pixels, Double Sided Strip detectors(Striplts).

Detector  DCH

Drift Chamber

40 layers of ≈ 1cm cells parralel to beam line. Provide momentum and \frac{dE}{dx} for low momentum particles(p < 700MeV ). ≈ 10000 channels Ocuupancy R& D: Geometry Gas mixture aaaa

Detector  DIRC

Detector of Internally Reflected Cherenkov light

Momentum range 0.7 − 4GeV Radiator: synthetic fused silica. Photon detectors outside field region. Radiatoin hard.

Detector  EMC and IFR

Electromagnetic and hadronic calorimeter

Electronamgnetic Calorimeter: Coverage 94%of 4Π CsI or LYSO cristals Crystal lenght 16 − 17.5X_{0} Radiatoin hard.

Instrumented Flux Return: Upgrade form TDC to BIRO Scintilators Iron reused from Babar SiPM

Physics Rare B Physics

B → τν

Precise SM prediction: Br(b → lν) = \frac{G^{2}_{F}m_{B}}{8Π}m_{l}^{2}(

Marcin Chrząszcz (IFJ PAN)  The SuperB factory  put date  13 / 13